OpenStudy (anonymous):

x + 15 + 2x= -5 -2x -25

OpenStudy (lgbasallote):

nice to know...

OpenStudy (anonymous):

Combine like terms. 15 + 3x = -30 - 2x Combine like terms again. Add 2x to both sides +2x +2x ------------------- 15 + 5x = -30 Now subtract 15 from both sides. -15 -15 ----------------------- 5x = -45 Divide. /5 /5x ----------- x = - 9 :) Hope this helps!

OpenStudy (anonymous):

Yes it did I'm checking my practice test to see if I'm doing it correctly. Thanks